Minutes of the 18th SOHO SWT Meeting

Goddard Space Flight Center, Greenbelt, MD 8 February 1996 Contents

1 Action Items SOHO SWT 18 3 2 Agree Agenda and Actions Revision 3 3 Status of S/C Commissioning 3 4 Status of Experiments 3 5 Status of Ground System 4 6 HOI Maneuver Plans 4 7 Priorities for the Coming Three Months 4 8 PR Matters 4 9 SOHO Science Report 5 10 Conferences 5 11 AOB 5 Annex 1: Agenda 6 Annex 2: List of Participants 8 Annex 3: SOHO Spacecraft Status and Plans 11 Annex 4: Roll Off-Pointing Anomalies 26 Annex 5: Status of Experiments 38

GOLF (A. Gabriel) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 VIRGO (C. Fröhlich) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 SOI-MDI (P. Scherrer) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 SUMER (P. Lemaire) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 CDS (R. Harrison) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 LASCO (G. Brueckner) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 SWAN (J.-L. Bertaux) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 CELIAS (D. Hovestadt) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 COSTEP (H. Kunow) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 ERNE (J. Torsti) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

Annex 6: SOHO Ground Operations Status 99 Annex 7: SOHO HOI Maneuver Planning 116 Annex 8: CDS Science Operations Memo #1 124 Annex 9: Summary of PRWG Meeting 131 Annex 10: Draft Layout of SOHO-News Press Releases 135 Annex 11: SOHO Science Related Conferences 138

2

1 Action Items SOHO SWT 18

Action 18-1: on PIs to check data quality and completeness of CD-ROMs. Input to PS by 28 February.

Action 18-2: on PIs to provide additional names for distribution list of PR package which will be circulated prior to the 2 May media event. Input to PS before 1 April.

Action 18-3: on PIs to provide input for the next SOHO Science Report. Input to PS before 15 March.

2 Agree Agenda and Actions Revision

Agenda — annex 1 List of Participants — annex 2

Actions revision

Action 17-1: closed on PS to clarify issue of experiment movements during PTCS calibration.

Action 17-2: closed on PIs to provide power-off recovery procedures to PS before Dec 15.

Action 17-3: closed on PS to clarify with FOT execution of emergency switch-off procedures on direct PI request.

Action 17-4: closed on PIs to provide e-mail addresses of SPWG members. Input to B.Fleck before Dec 15.

Action 17-5: closed on PS to contact R.Harrison and arrange free block for SOHO presentations at EGS (Dec 15).

Action 17-6: closed on PS to arrange special SOHO session at 1996 SPD meeting (Dec 15).

Action 17-7: open (except SUMER and ERNE) on PIs to provide bibliography information (list of papers with abstracts, database used). Input to B.Fleck before January 30.

Action 17-8: closed on PIs to provide info on space requests (EOF + EAF) to A.Poland (Dec 15).

3 Status of S/C Commissioning

F. Vandenbussche gave a presentation to D/Sci on Wednesday morning, Feb 7 (for handout see annex 3). M. Hervieux gave a presentation on the roll off-pointing anomalies (see annex 4).

4 Status of Experiments

– see annex 5 There is an action on the PIs to check the data quality and completeness of CD-ROMs and report to PS before 28 February (Action 18-1).

5 Status of Ground System

K. Walyus gave a presentation to D/Sci on Wednesday morning, Feb 7 (for handout see annex 6).

6 HOI Maneuver Plans

B. Short gave a presentation on the HOI maneuver scheduled for February 14 (see annex 7).

7 Priorities for the Coming Three Months

Priorities as presented by Richard Harrison, discussed and agreed upon by SWT:

1. Finish commissioning and get into science operations (individual)

2. Each group produce publicity material for 2 May event (may need dedicated ops/target)

3. Start monitoring & calibration operations/observations, including intercalibration sequences

4. Quick return & first look observations

5. Start some basic JOPs on 1 April

6. Synoptic observations

The “first month of SOHO science” starts on 1 April. For CDS plans for the first month see annex 8 (CDS Science Operations Memo #1).

Priorities for the first 3 months of SOHO science are:

1. Coronal Holes

2. Streamers

3. Diffuse Corona

4. Quiet Network

SOWG to prioritize JOPs in compliance with these 4 topics. CME studies may get some emphasis towards the end of the first 3 months.

Two months period of continuous DSN contact: not yet clear when DSN will be able to support; request submitted to have it as early as possible but no later than July/August timeframe.

8 PR Matters

For a summary of the PRWG meeting which took place on Tuesday, February 6, in preparation of the 2 May media event see annex 9. Tentative presenters for the three topics at the two parallel events in ESA HQ in Paris and in GSFC are:

i) Solar Interior: A. Gabriel, P. Scherrer ii) Solar Atmosphere: R. Harrison, G. Brueckner iii) Solar Wind and Heliosphere: J.-L. Bertaux, P. Bochsler It was proposed to have M. Huber and G. Withbroe as moderator for this media event.

It was emphasized that there should be one (1) joint ESA/NASA press kit distributed before the 2 May event, rather than an ESA kit in Europe and a NASA one in the US. PIs to provide additional names/addresses for the distribution list of this press kit (Action 18-2).

V. Domingo presented a draft layout for the SOHO-NEWS press releases (see annex 10). Comments and suggestions are welcome. The final style will be chosen such that the press releases can be easily made available as PostScript files.

V. Domingo will arrange for distribution of Nigel Calder’s book “Beyond this World” (ESA BR-112) to PIs.

9 SOHO Science Report

ESA’s Publications Division will produce and circulate future SOHO Science Reports. The next SOHO Science Report (the first in the new style) should be ready by mid of April, to be circulated together with the PR media package in preparation of the 2 May event. There is an action on the PIs to provide their input for the next Science Report before March 15 (Action 18-3). Color diagrams are welcome. The PIs are also requested to provide names/addresses for an extended distribution list of the SOHO Science Report. The present distribution list includes the PIs, all Co-Is, ESA+NASA management, 65 solar ground-based observatories around the world plus some individuals who expressed their interest.

10 Conferences

For a complete list of SOHO science related conferences see annex 11. Special SOHO sessions have been arranged for at:

– EGS in The Hague, NL (May 6-10)

– AGU in Baltimore, MD (May 20-24) (A. Poland urgently needs input from PIs!)

– SPD in Madison, WI (June 8-13) (organizers: J. Gurman and B. Fleck)

– COSPAR in Birmingham, UK (June 15-20)

IAU XXIIIrd General Assembly (18-30 August 1997 in Kyoto, Japan): There are plans to organize a Seismology Symposium by Comm. 12 and a Joint Discussion (1 1/2 days) on the coronal aspects by Comm. 10. B. Fleck to contact F.-L. Deubner (Pres. Comm. 12) and O. Engvold (Pres. Comm. 10) for more details.

11 AOB

Next Meeting From now on the SPWG meetings will be held on the 3rd Friday each month. Next SWT meeting: 16/17 May 1996 (again in conjunction with a SPWG meeting)

Annex 1: Agenda

6

Annex 2: List of Participants

8

Annex 3: SOHO Spacecraft Status and Plans

11

Annex 4: Roll Off-Pointing Anomalies

26

Annex 5: Status of Experiments

GOLF (A. Gabriel)

VIRGO (C. Fröhlich)

41

SOI-MDI (P. Scherrer)

45

SUMER (P. Lemaire)

48

CDS (R. Harrison)

55

LASCO (G. Brueckner)

65

SWAN (J.-L. Bertaux)

68

CELIAS (D. Hovestadt)

70

COSTEP (H. Kunow)

78

ERNE (J. Torsti)

86

Annex 6: SOHO Ground Operations Status

99

Annex 7: SOHO HOI Maneuver Planning

116

Annex 8: CDS Science Operations Memo #1

124

Annex 9: Summary of PRWG Meeting

131

Annex 10: Draft Layout of SOHO-News Press Releases

135

Annex 11: SOHO Science Related Conferences

138